# Accurate Unsupervised Photon Counting from Transition Edge Sensor Signals

### v2

@@ IV.1 Validation
Before looking at performance metrics, a sanity check is done to validate the basic characteristics of the Synthetic Uniform dataset. This is done for the data from the different coherent sources (all with different mean photon numbers). Since coherent sources are used to generate the samples, a $g^{(2)}$ (second-order correlation) of 1 is expected. This quantity is defined in terms of the first two moments of the photon number distribution, as
$g^{(2)}=\frac{\langle n^{2}\rangle-\langle n\rangle}{\langle n\rangle^{2}}.$ (27)
We use the $g^{(2)}$ as a validation metric both ways by making sure the statistics of the light are correct and that the generated statistics using the numerical methods follow the physics of the system.
Figure 2: Computed second-order correlation for the different datasets (where markers are the mean photon number of the available coherent sources) and methods. In this figure, and the ones that follow, methods using a 1D latent space are represented by dotted lines, while those with 2D latent spaces are shown with solid lines.
In Fig. 2 we can see that every method has a $g^{(2)}$ close to 1 for most datasets. All methods consistently get farther from one as the mean photon number increases, the lack of resolution for high photon numbers explains this behaviour. Additionally, the number of signals associated with the high mean is limited compared to the low mean cases. The lack of resolution is especially present for the method based on the maximum value of the signals, since it cannot resolve photon numbers higher than 10 in our dataset.
@@ IV.2 Confidence
Figure 3: Confidence of photon number clusters for the different methods using the Synthetic Uniform dataset. In this figure, and the ones that follow, methods using a 1D latent space are represented by dotted lines, while those with 2D latent spaces are shown with solid lines.
Figure 4: Confidence of photon number clusters for the different methods using the Synthetic Geometric dataset
The number of clusters considered in the confidence plots is defined using the AIC and BIC information criteria and other considerations. First, the last cluster is always removed since it often offers an artificially high confidence considering there is no other cluster to overlap with farther in the latent space. Additionally, regions associated with multiple photons described by a uniform density are ignored. This is done since regions of uniform density can be described by an arbitrarily large number of Gaussians.
@@ V.1 Qualitative Analysis
Figure 5: Kernel density estimation of the low dimensional embedding of TES signals generated by (5(a)) PCA 2D, (5(b)) t-SNE 2D, (5(c)) UMAP 2D, (5(d)) PCA 1D, (5(e)) t-SNE 1D, (5(f)) UMAP 1D.
@@
PCA is the first interesting method, since it was previously used for this task. We observe clear clusters, and the samples followed the expected arc-like structure presented in Fig. 5(a) and observed in previous work [35].
@@ V.2 Limits for Parametric Implementations
The training process to generate a network with the reported performance for the Synthetic Uniform and Geometric datasets required a fair amount of tuning to give satisfying results, which is not ideal for experimental setups. We mainly attribute this problem to the limited amount of training data, which makes it easy to overfit the model to the training data. More precisely, by learning local data structures the neural network learns less generalized features which limits its capacity to make predictions. This family of neural networks is therefore more reliant on having access to a large training dataset, since it needs examples for a wider range of fine signal features. This limits the performance capabilities demonstrated in this work, however, with a larger training set the neural networks can have prediction capabilities similar to the transformation of their non-parametric implementation. To verify this intuition, we used the Synthetic Large dataset previously mentioned in section III.7. Using the $u=300000$ signals, we trained a small feedforward neural network (5 linear layers of size 300). We present in Fig. 6 that with sufficient data, this network offers advantageous confidence values compared to previously used techniques in one-dimension.
@@ V.3 Impact of Embedding Dimension
The analysis of the dimensionality reduction techniques in this work assumes that the underlying true classes are associated with the photon numbers. This gives satisfying results because the traces for each photon number follow a clear pattern that different methods can easily capture. However, additional considerations are needed to solve the photon number classification problem. First, cluster distinguishability inside the low-dimensional representations is possible because the underlying structures of photon numbers are dominant in comparison to other characteristics like noise. Additionally, the dimensionality reduction techniques are only aware of data structure at different scales and never explicitly have a grasp of the physical system. We emphasize this property since it makes the method almost completely independent of the statistics of the measured light and does not require prior knowledge of the light source. To come back to the data structures, when methods encode data in a low dimensional space they need to find a representation that describes the entire complexity of the signals. This means that noise and photon number structures are equally preserved in the embedding. If enough noise structures exist, the method will not have enough space in a single dimension to represent this variety, and the resulting embedding can show excessive broadening of clusters. The constraint of preserving structures in the data limits the potential of finding well separated clusters in lower dimensional embeddings. This is a reason why it is easier to find an embedding with well-separated clusters in two-dimensional spaces, even if the underlying classes we wish to identify are contained in a single dimension: the photon number.
@@ V.5 Outlier Detection
A one-dimensional embedding is efficient from a computational point of view, since the clustering problem can be translated into a sorted array search. However, depending on the use case, we argue that two dimensions may offer deeper insight due to their capacity to capture a wider range of structures. For example, if temporally uncorrelated light overlaps with the light modes one seeks to analyse, then a single dimension is likely not enough space to correctly capture the photon-number statistics of the modes under analysis. Adding to what is mentioned in the previous section, the noise becomes an additional structure to represent, and effectively the proportion of information that the method can allocate to the photon number structure is reduced. This is shown in Fig. 7(a) where we use the Noise dataset (section III.7) and observe cluster broadening due to the presence of temporally uncorrelated photons. In this case, the two-dimensional representation becomes more useful, cf. Fig. 7(b), to describe the complexity of the dataset.
@@
Figure 8: In the centre, we present a low dimensional representation using UMAP of a dataset containing signals from a system with temporally uncorrelated noise. Each cluster in the kernel density estimation is identified using lower case letters, and each graph, identified using the associated upper case letter, represents the signals in each labelled clusters. (8a), (8d), (8g), and (8h) give the temporally correlated photon numbers 0 to 3. (8b) and (8c) are associated to uncorrelated signals, with zero photons correlated before and after the trigger time. (8e) and (8f) are single photons at the trigger time and uncorrelated signals before and after the trigger.
@@
We also noticed that methods that preserve local structures tend to create clearer clusters for noise structures, facilitating the clustering task. This effect is seen in Fig. 8 where the uncorrelated noise is found on curve structures and photon numbers in tear-like shapes. If we look closely at the content of these clusters, we see that it is possible to identify signals of uncorrelated single photons before the trigger time (cluster 8c) and after the trigger time (cluster 8b). Similarly, we find uncorrelated single photons combined with correlated single photons in clusters 8e and 8f. In clusters 8a, 8d, 8g, and 8h we find the standard photon numbers 0 to 3 without uncorrelated light. Similar analysis could be done using more traditional methods like PCA, however the clustering becomes significantly harder. This lack of cluster structure is visually demonstrated in Fig. 7 where the uncorrelated light becomes a broadening of the temporally correlated photon numbers.
@@
We note that the Gaussian Mixture Model is not as effective in clustering noise features, especially considering a photon number embedding from UMAP. We found that methods like HDBSCAN, which is a hierarchical density-based clustering technique, are well-suited for UMAP embedding [57]. This technique has the main advantage of working on clusters that do not follow a Gaussian structure, which is adequate for noise clusters that can have a variety of shapes.
V.6 Impact of Gaussian Mixture Model
@@
We note that one-dimensional results for t-SNE offer clusters that follow top-hat-like distributions, cf. Fig. 5. This feature decreases the confidence results, but not the actual potential clustering over this embedding. For a more accurate representation of t-SNE clusters, we use a generalized Gaussian distribution to represent the probability density of each cluster defined as
@@
In these equations, $\mu_{n}$, $\sigma_{n}^{2}$, and $\Gamma$ are respectively the mean and variance of a given photon number cluster and the Gamma function. In Fig. 9(a) we present a qualitative representation of the fit quality of t-SNE embedding using the standard and generalized Gaussians functions.
@@
V.7 Potential Applications
Based on the benchmarks, the dimensionality reduction techniques that focus on local structure preservation offer the best low-dimensional representation of the transition edge sensor signals.
@@ V.7 Potential Applications
Considering the previous performance results, neural networks (5 linear layers of 300 neurons) can offer a trustworthy and interpretable low-dimensional representation of the TES traces. The condition necessary for this method to be accurate is to provide a balanced dataset containing the range of photon numbers we want to detect. It is essential to understand that the network cannot predict photon number outside the trained range, since it never learned an embedding for these signals. The training data restricts the learned transformation. Moreover, our results suggest that a small neural network implemented in a Field Programmable Gate Array (FPGA) [59] could replace currently used methods like trace area and PCA [25] to process the TES traces directly. With this type of hardware we believe real-time processing can be achieved considering TESs have a dead time of a few microseconds and knowing our CPU implementation can process a TES signal of 200 points in $4.9\mu s$. This value is obtained using a laptop with a clockspeed of 3.2 GHz, 8 cores and 16 threads.
We emphasized using a close-to-uniform distribution to train the network, since it becomes equally optimized for every class (photon number). Following the example used to benchmark the different methods, the use of a coherent source with tunable mean photon number is more than sufficient to create a balanced dataset. It is therefore possible to create suitable conditions only using a laser and tunable attenuation. We could also imagine using a high mean photon number thermal source, depending on the available equipment.
@@
V.8 Future work
While testing the different methods, the clustering step (Gaussian Mixture Model) was particularly sensitive to the initialization process. Often some manual adjustments had to be done to guarantee the quality of the results. To further improve the quality and robustness of photon number classification, future work could explore clustering techniques that may be better aligned with the novel methods introduced in this study. This way it could be possible to completely automate the photon number classification process even for low visibility clusters.
Accessing the ground truth in the case of photon number classification remains difficult, and further validation would be desirable to guarantee the performance of the proposed methods. We propose using the joint probability distribution of photon pairs to benchmark dimensionality reduction methods. In more detail, in a perfect system, photon pairs should have the same photon number in both modes. Experimentally, loss and misassigned photons broaden the joint probability distribution, which would be otherwise perfectly diagonal [61]. With this in mind, we expect broadening effects associated with the numerical analysis. This way, the width of the joint probability distribution becomes an experimental tool to quantify the performance of numerical techniques.

### v3

@@ IV.1 Validation
Before looking at performance metrics, a sanity check is done to validate the basic characteristics of the Synthetic Uniform dataset. This is done for the data from the different coherent sources (all with different mean photon numbers). Since coherent sources are used to generate the samples, a $g^{(2)}(0)$ (second-order coherence) of 1 is expected. This quantity is defined in terms of the first two moments of the photon-number distribution, as
$g^{(2)}(0)=\frac{\langle n^{2}\rangle-\langle n\rangle}{\langle n\rangle^{2}}.$ (27)
We use the $g^{(2)}(0)$ as a validation metric both ways, by making sure the statistics of the light are correct and that the generated statistics using the numerical methods follow the physics of the system.
Figure 4: Computed second-order coherence for each method used on the individual datasets composing the Synthetic Uniform dataset (where markers are the mean photon number of the available coherent sources). In this figure, and the ones that follow, methods using a 1D latent space are represented by dotted lines, while those with 2D latent spaces are shown with solid lines.
In Fig. 4 we can see that every method has a $g^{(2)}(0)$ close to 1 for most datasets. All methods consistently get farther from one as the mean photon number increases, the lack of resolution for high photon numbers explains this behaviour. Additionally, the number of signals associated with the high mean is limited compared to the low mean cases. The lack of resolution is especially present for the method based on the maximum value of the signals, since it cannot resolve photon numbers higher than 10 in our dataset.
@@ IV.2 Confidence
Figure 5: Confidence of photon-number clusters for the different methods using the Synthetic Uniform dataset. In this figure, and the ones that follow, methods using a 1D latent space are represented by dotted lines, while those with 2D latent spaces are shown with solid lines.
Figure 6: Confidence of photon-number clusters for the different methods using the Synthetic Geometric dataset
The number of clusters considered in the confidence plots is defined using the AIC and BIC information criteria and other considerations. First, the last cluster is always removed since it often offers an artificially high confidence considering there is no other cluster to overlap with further in the latent space. Additionally, regions associated with multiple photons described by a uniform density are ignored. This is done since regions of uniform density can be described by an arbitrarily large number of Gaussians.
@@ V.1 Qualitative Analysis
Figure 7: Kernel density estimation of the low dimensional embedding of TES signals generated by (7(a)) PCA 2D, (7(b)) t-SNE 2D, (7(c)) UMAP 2D, (7(d)) PCA 1D, (7(e)) t-SNE 1D, (7(f)) UMAP 1D. The clusters are labelled by sorting the clusters by the mean area of the samples inside each cluster, since the trace area is expected to increase with the photon number.
@@
PCA is the first interesting method, since it was previously used for this task. We observe clear clusters, and the samples follow the expected arc-like structure presented in Fig. 7(a) and observed in previous work [37].
@@ V.2 Limits for Parametric Implementations
The training process to generate a network with the reported performance for the Synthetic Uniform and Geometric datasets required a fair amount of tuning to give satisfying results, which is not ideal for experimental setups. We mainly attribute this problem to the limited amount of training data, which makes it easy to overfit the model to the training data. More precisely, by learning local data structures the neural network learns less generalized features which limits its capacity to make predictions. This family of neural networks is therefore more reliant on having access to a large training dataset, since it needs examples for a wider range of fine signal features. This limits the performance capabilities demonstrated in this work, however, with a larger training set the neural networks can have prediction capabilities similar to the transformation of their non-parametric implementation. To verify this intuition, we used the Synthetic Large dataset previously mentioned in section III.7. Using the u=300 000u=300\,000 signals, we trained a small feedforward neural network (5 linear layers of size 300). We present in Fig. 8 that with sufficient data, this network offers advantageous confidence values compared to previously used techniques in one-dimension, which was not the case using less data. We limit this comparison to one-dimensional embeddings, as our focus is on techniques suitable for real-time signal processing, and the labelling task remains computationally efficient only in one dimension.
@@ V.3 Impact of Embedding Dimension
The analysis of the dimensionality reduction techniques in this work assumes that the underlying true classes are associated with the photon numbers. This gives satisfying results because the traces for each photon number follow a clear pattern that different methods can easily capture. However, additional considerations are needed to solve the photon-number classification problem. First, cluster distinguishability inside the low-dimensional representations is possible because the underlying structures of photon numbers are dominant in comparison to other characteristics like noise. Additionally, the dimensionality reduction techniques are only aware of data structure at different scales and never explicitly have a grasp of the physical system. We emphasize this property since it makes the method almost completely independent of the statistics of the measured light and does not require prior knowledge of the light source. To come back to the data structures, when methods encode data in a low dimensional space they need to find a representation that describes the entire complexity of the signals. This means that noise and photon-number structures are equally preserved in the embedding. If enough noise structures exist, the method will not have enough space in a single dimension to represent this variety, and the resulting embedding can show excessive broadening of clusters. The constraint of preserving structures in the data limits the potential of finding well separated clusters in lower dimensional embeddings. This is a reason why it is easier to find an embedding with well-separated clusters in two-dimensional spaces, even if the underlying classes we wish to identify are contained in a single dimension: the photon number.
@@ V.5 Outlier Detection
A one-dimensional embedding is efficient from a computational point of view, since the clustering problem can be translated into a sorted array search. However, depending on the use case, we argue that two dimensions may offer deeper insight due to their capacity to capture a wider range of structures. For example, if temporally uncorrelated light overlaps with the light modes one seeks to analyze, then a single dimension is likely not enough space to correctly capture the photon-number statistics of the modes under analysis. Adding to what is mentioned in the previous section, the noise becomes an additional structure to represent, and effectively the proportion of information that the method can allocate to the photon-number structure is reduced. This is shown in Fig. 9(a) where we use a single channel in the Noise dataset (section III.7) and observe cluster broadening due to the presence of temporally uncorrelated photons. In this case, the two-dimensional representation becomes more useful to describe the complexity of the dataset (see Fig. 9(b)).
@@
Figure 10: In the centre, we present a low dimensional representation using UMAP of a dataset containing signals from a system with temporally uncorrelated noise. Each cluster in the kernel density estimation is identified using lower case letters, and each graph, identified using the associated upper case letter, represents the signals in each labelled clusters. (a), (d), (g), and (h) give the temporally correlated photon numbers 0 to 3. (b) and (c) are associated with uncorrelated signals, with zero photons correlated before and after the trigger time. (e) and (f) are single photons at the trigger time and uncorrelated signals after and before the trigger. The remaining samples in cluster (i) are associated with the higher photon numbers and remaining noise photons.
@@
We also noticed that methods that preserve local structures tend to create clearer clusters for noise structures, facilitating the clustering task. This effect is seen in Fig. 10 where the uncorrelated noise is found on curve structures and photon numbers in circular shapes. If we look closely at the content of these clusters, we see that it is possible to identify signals of uncorrelated single photons before the trigger time (cluster 10b) and after the trigger time (cluster 10c). Similarly, we find uncorrelated single photons combined with correlated single photons in clusters 10e and 10f. In clusters 10a, 10d, 10g, and 10h we find the standard photon numbers 0 to 3 without uncorrelated light. Similar analysis could be done using more traditional methods like PCA, however the clustering becomes significantly harder. This lack of cluster structure is visually demonstrated in Fig. 9 where the uncorrelated light becomes a broadening of the temporally correlated photon numbers.
The remaining samples in cluster 10i are associated with the higher photon numbers and remaining noise photons. Since a quasi-thermal light source is used, these signals only represent a small portion of the dataset.
@@
We note that the Gaussian Mixture Model is not as effective in clustering noise features, especially considering a photon-number embedding from UMAP. We found that methods like HDBSCAN, which is a hierarchical density-based clustering technique, are well-suited for UMAP embedding [60]. This technique has the main advantage of working on clusters that do not follow a Gaussian structure, which is adequate for noise clusters that can have a variety of shapes.
V.6 Correlated Photons
To further quantify the impact of noise photons and demonstrate that the clusters identified by UMAP correspond to real physical features, we examine how photon-number assignment influences the reconstructed photon-number distributions of the correlated channels in the Noise dataset. In this system, we expect thermal photon statistics on each individual channel, with second-order coherence approaching $g^{(2)}(0)=2$, along with an added noise contribution from pump leakage. The resulting statistics should therefore lie between Poissonian and thermal behaviour, yielding a $g^{(2)}(0)$ value between 1 and 2, where 1 corresponds to purely Poissonian light and 2 to purely thermal light. As a baseline, we begin by analyzing data obtained when we optimize the pump filtering, i.e. minimize the amount of temporally uncorrelated noise, we obtain values of $g^{(2)}(0)$ of approximately $1.848\pm 0.008$ for the first channel and $1.879\pm 0.004$ for the second. The uncertainties reflect the spread in values obtained using different numerical methods. Given the low mean photon number ($\bar{n}\sim 1.01$), the computed values provide a reliable approximation of the true $g^{(2)}(0)$ parameter for a thermal light source, largely independent of the specific numerical method employed.
Next, we analyze data obtained with increased noise by reducing the pump filtering (which is the data found in the Noise dataset). Using one-dimensional PCA, we obtain $g^{(2)}(0)$ values of 1.514 and 1.592 for the two channels, respectively. A comparable result is observed with the integrated area method, yielding values of 1.431 and 1.459. However, upon using the extra features obtained by UMAP we can reject events associated with noise and thus correct some potential dark counts. As a result, the $g^{(2)}(0)$ values increase to 1.695 and 1.699, bringing them closer to the expected thermal regime. This emphasizes the effectiveness of UMAP-based preprocessing in isolating and mitigating noise.
In parallel, we estimate the noise reduction factor ($\mathrm{NRF}$) of the clean source, defined as:
$\mathrm{NRF}=\frac{\mathrm{Var}(n_{1}-n_{2})}{\langle n_{1}+n_{2}\rangle},$ (28)
where $n_{1}$ and $n_{2}$ denote the photon numbers in the two channels [61]. For comparison, a Poissonian source yields $\mathrm{NRF}=1$, whereas twin beams states exhibit $\mathrm{NRF}<1$, with values increasing in the presence of loss or noise. Across our numerical methods (area, PCA, and UMAP), we compute a baseline value of $\mathrm{NRF}=0.772\pm 0.005$.
For the $\mathrm{NRF}$, we compute values of 0.826 (area), 0.810 (PCA), and 0.811 (UMAP), showing only a modest increase from the baseline. Notably, the $\mathrm{NRF}$ shows limited sensitivity to the choice of numerical method. We attribute this to the low system efficiency, characterized by a transmission coefficient of approximately $\eta\approx 22\%$. Since the $\mathrm{NRF}$ scales as $1-\eta$, it approaches 1 in low-efficiency regimes, where it becomes less sensitive to noise. This contrasts with $g^{(2)}(0)$, which remains loss-insensitive. A simulation illustrating this effect is available in a public repository [54].
To further illustrate the effect of photon-number misclassification, Fig. 11 presents the scaled difference between the joint photon-number distributions derived from PCA and UMAP. The most prominent deviation stems from PCA misclassifying 0-photon events approximately 4% of the time due to noise influence, reinforcing the advantage of UMAP for denoising and accurate state reconstruction. This effect is also reflected in the predicted mean photon number across the different techniques, as temporally uncorrelated noise primarily decreases the number of predicted 0-photon events. We observe a decreasing trend in the estimated mean photon number that correlates with each method’s ability to correctly identify and reject noise. The area method yields a mean photon number of 0.697, as it effectively integrates the total signal over the full time interval. PCA, which partially filters traces based on temporal features, reduces the mean to 0.637. UMAP, with its stronger noise discrimination capabilities, further decreases the estimated mean to 0.566.
Figure 11: Scaled difference between the joint photon-number distributions computed using UMAP (2D) and PCA (1D) on the Noise dataset (light source described in Sec. III.7). For reference, the unscaled joint distribution is shown for the first few photon-number states.
V.7 Impact of Gaussian Mixture Model
@@
One-dimensional results for t-SNE offer clusters that follow top-hat-like distributions, cf. Fig. 7. This feature decreases the confidence results, but not the actual potential clustering over this embedding. For a more accurate representation of t-SNE clusters, we use a generalized Gaussian distribution to represent the probability density of each cluster defined as
@@
In these equations, $\mu_{n}$, $\sigma_{n}^{2}$, and $\Gamma$ are respectively the mean and variance of a given photon-number cluster and the Gamma function. In Fig. 12(a) we present a qualitative representation of the fit quality of t-SNE embedding using the standard and generalized Gaussian functions.
@@
V.8 Potential Implementations
Based on the benchmarks, the dimensionality reduction techniques that focus on local structure preservation offer the best low-dimensional representation of the transition-edge sensor signals.
@@ V.8 Potential Implementations
Considering the previous performance results, neural networks (5 linear layers of 300 neurons) can offer a trustworthy and interpretable low-dimensional representation of the TES traces. The condition necessary for this method to be accurate is to provide a balanced dataset containing the range of photon numbers we want to detect. It is essential to understand that the network cannot predict photon numbers outside the trained range, since it never learned an embedding for these signals. The training data restricts the learned transformation. Moreover, our results suggest that a small neural network implemented in a Field Programmable Gate Array (FPGA) [63] could replace currently used methods like trace area and PCA [27] to process the TES traces directly. With this type of hardware we believe real-time processing can be achieved considering TESs have a dead time of a few microseconds and knowing our CPU implementation can process a TES signal of 200 points in $4.9\mu s$. This value is obtained using a laptop with a clock speed of 3.2 GHz, 8 cores and 16 threads.
We emphasized using a close-to-uniform distribution to train the network, since it becomes equally optimized for every class (photon number). Following the example used to benchmark the different methods, the use of a coherent source with a tuneable mean photon number is more than sufficient to create a balanced dataset. It is therefore possible to create suitable conditions only using a laser and tuneable attenuation. We could also imagine using a high mean photon number thermal source, depending on the available equipment.
V.9 Future Work
While testing the different methods, the clustering step (Gaussian Mixture Model) was particularly sensitive to the initialization process. Often some manual adjustments had to be done to guarantee the quality of the results. To further improve the quality and robustness of photon-number classification, future work could explore clustering techniques that may be better aligned with the novel methods introduced in this study. This way it could be possible to completely automate the photon-number classification process even for low visibility clusters.
@@ V.9 Future Work
Accessing the ground truth for photon-number classification remains a significant challenge, and further validation is needed to robustly assess the performance of the proposed methods. To this end, we leverage the joint probability distribution of twin beams to evaluate the impact of dimensionality reduction techniques (see Sec. V.5). This distribution serves as an intuitive diagnostic tool, analogous to a confusion matrix commonly used in classification tasks. In our current implementation, generating twin beams with high mean photon numbers and strongly correlated channels is not feasible. Nevertheless, the distribution proves useful for assessing the filtering capabilities of the methods under study. A future experiment with higher photon numbers, lower loss, and reduced noise would offer deeper insight into the potential of this numerical processing. Given these considerations, we anticipate broadening effects introduced by the numerical techniques; thus, the width of the joint probability distribution becomes a practical experimental observable to quantify their performance.